OpenStudy (jazilove):

I WILL GIVE MEDAL ABC has vertices A(0, 6), B(4, 6), and C(1, 3). Sketch a graph of ABC and use it to find the orthocenter, including any necessary points or slopes you had to derive. (PLEASE PLEASE PLEASE help i need help fast)

OpenStudy (wolf1728):

JaziLove Do you know how to graph it?

OpenStudy (jazilove):

Not really

OpenStudy (jazilove):

Well, i think so,, but i dont know what to do after that

OpenStudy (victoriasushchik):

The point where the three altitudes of a triangle intersect. The altitude of a triangle (in the sense it used here) is a line which passes through a vertex of the triangle and is perpendicular to the opposite side. There are therefore three altitudes possible, one from each vertex Circumcentre Definition: Located at intersection of the perpendicular bisectors of the sides. The circumcenter is also the center of the triangle's circumcircle - the circle that passes through all three of the triangle's vertices. Incentre Definition: Located at intersection of the angle bisectors. The incenter is the center of the triangle's incircle, the largest circle that will fit inside the triangle and touch all three sides Try drawing these triangles on graph paper. You will see these three centres.

OpenStudy (jazilove):

I'm still confused on how to work the problem

OpenStudy (jazilove):

@pooja195 please help

OpenStudy (jazilove):

anyone please help

OpenStudy (wolf1728):

I'm still here - just finishing up a graphic

OpenStudy (jazilove):

oh ok

OpenStudy (jazilove):

hello

OpenStudy (jazilove):

@wolf1728

OpenStudy (wolf1728):

Okay, I'm back. I had to redraw the graphic to include the "6"

OpenStudy (wolf1728):

In the graph the "x" axis has numbers that go left to right the "y" axis has numbers that go from bottom to top

OpenStudy (jazilove):

how would i solve it step by step

OpenStudy (wolf1728):

So, to plot point A, you would need to put a point at x=0 and y=6 I just put it into the graphic

OpenStudy (wolf1728):

I just plotted the points for B and C and put those in the graphic

OpenStudy (jazilove):

ok

OpenStudy (jazilove):

is there more?

OpenStudy (wolf1728):

Yes because now you have to calculate where the orthocenter is (wow this problem requires a lot doesn't it?)

OpenStudy (jazilove):

yes it does *sigh*

OpenStudy (jazilove):

how do i fond the orthocenter

OpenStudy (wolf1728):

I just drew a few lines to make a triangle of the 3 points. I think I might do a quick web search to see about "orthocenter"

OpenStudy (jazilove):

alright

OpenStudy (wolf1728):

GEEZ - here's a site that tells how to do it: https://www.easycalculation.com/analytical/orthocenter-triangle.php Seems you have to find the slopes of the lines (Guess I'll read some more)

OpenStudy (jazilove):

whoa, this is a very difficult problem, thank you for helping for this long

OpenStudy (wolf1728):

Okay I'll keep going if you want

OpenStudy (jazilove):

please

OpenStudy (wolf1728):

Sure

OpenStudy (wolf1728):

Slope of line AB =0 Slope of line AC = (6-3)/(0-1) = 3/-1 = -3 Slope of line CB = (3-6) / (1-4) = -3/-3 = 1 Now it looks as if we move on to Step 2 on the orthocenter page

OpenStudy (jazilove):

looks like itxD

OpenStudy (wolf1728):

Slope of AD= 1/(slope BC) = 1/1 = 1 slope of BE = 1/(slope AC) = (1/-3) slope of CF = 1/(slope AB) = (1/0) = Infinite

OpenStudy (jazilove):

im kind of confused about the last part

OpenStudy (wolf1728):

Heck, I have found this exact problem worked out here: https://www.wyzant.com/resources/answers/2717/abc_has_vertices_a_0_6_b_4_6_and_c_1_3_find_the_orthocenter_of_abc_list_your_steps Incidntally, the anwer turns out to be (1,5) or x=1 y=5 which is right where I drew it !!

OpenStudy (jazilove):

wow haha thanks for everything

OpenStudy (wolf1728):

Okay you are welcome I really didn't show you all the equations to get the answer but I think that site shows you how to do it. Here's my final graphic with the orthocenter drawn in!

OpenStudy (wolf1728):

Thanks for the medal :-)